OpenStudy (anonymous):

Ricardo Gutierrez has a box of coins that he uses when playing poker with friends. The box contains 28 coins, consisting of pennies, dimes, and quarters. The number of pennies equal to the number of dimes, and the total value is $2.34. How many of each denomination of coins does he have?

OpenStudy (anonymous):

lets call the number of pennies \(x\) then the number of dimes is also \(x\) since there are 28 coins, the number of quarters must therefore be \(28-2x\) total amount of money in pennies is \(234\) and that must be \(x+10x+25(28-2x)\) so solve \[x+10x+25(28-2x)=234\] for \(x\)

OpenStudy (anonymous):

-39x+466 ?

OpenStudy (anonymous):

????

hartnn (hartnn):

that is correct left side, so -39x+466=0 then x= 466/39 = ??

OpenStudy (anonymous):

11.?

OpenStudy (anonymous):

0? @hartnn?

hartnn (hartnn):

11.949 or 12 approximately.

OpenStudy (anonymous):

So there were 11.949 quarters and 234 pennies and dimes?

OpenStudy (anonymous):

@hartnn ?

hartnn (hartnn):

no,the dimes and pennies are of same number = 12

OpenStudy (anonymous):

and the quarters are?

hartnn (hartnn):

28-2x=28-2*12=28-24=4

OpenStudy (anonymous):

So 11.949 penies and dimes and 4 quarters?

hartnn (hartnn):

yes,to be exact.

OpenStudy (anonymous):

Find a value for x so that 4/10=-2/x?

OpenStudy (anonymous):

will it be x=5?

hartnn (hartnn):

u forgot negative sign, x=-5

OpenStudy (anonymous):

oh, i sure did forget the negatvie sign lol, thanks

OpenStudy (anonymous):

What is the converse and inverse of the statement? If charles wrote a sonata, then snoppy did not dance a jig.